True or False? Nigerias currency is the naira (NGN); Malawis currency is the kwacha (MWK). The annual rate of appreciation
Answer #1

Let the exchange rate at the time of investment be x NGN/MWK be x

Purchase value in terms of naira=100000*x NGN

Exchange rate after 1 year=x*(1-4%)=0.96 NGN/MWK

Sale value in terms of naira=106000*0.96x NGN

Nominal rate of return=(106000*0.96x-100000x)/(100000x)=1.76%

So, Given statement is false